OpenStudy (hba):

A particle moves along x-axis.It starts from rest at origin O and its acceleration after t sec is given by a=3-t . Calculate the displacment from O when it is at rest.

OpenStudy (hba):

@saifoo.khan

OpenStudy (hba):

Can you do it ?

OpenStudy (jennychan12):

find antiderivative twice for a = 3-t

OpenStudy (saifoo.khan):

integrate. a=3-t v = 3t - t^2/2 + C Now set it equal to zero to get time. Then integrate again, find the disp.

OpenStudy (hba):

Therefore, the answer is 18. Something which appeared in the (NUST) test.

OpenStudy (hba):

@saifoo.khan

OpenStudy (saifoo.khan):

REALLY? O_O

OpenStudy (hba):

Yeah

OpenStudy (saifoo.khan):

An O'levels student who have Add math can solve this too! D: D: D:

OpenStudy (hba):

Hmm,good :)

OpenStudy (hba):

What about hyperbolic functions ? You've done that ?

OpenStudy (saifoo.khan):

give me the problem, then i can tlel.

OpenStudy (hba):

\[\int\limits_{}^{}e^x[\sinh^{-1} x+\frac{ 1 }{ \sqrt{1+x^2} }]dx=\]

OpenStudy (hba):

@saifoo.khan Like this.

OpenStudy (saifoo.khan):

INtegration by parts?

OpenStudy (saifoo.khan):

idk how to deal with sinh^-1

OpenStudy (hba):

So you don't know hyperbolic functions ?

OpenStudy (experimentx):

use e^x(f(x) + f'(x)) rule differentiation of arcsinh(x) = 1/sqrt(...)

OpenStudy (saifoo.khan):

I guess no, then. :S

OpenStudy (hba):

hmm No need to worry about that it's easy,you can do that later. What about finding the area from a diag ?

OpenStudy (saifoo.khan):

We have to learn that parallelogram thing. But i can find the area/vol of a curve and stuff

OpenStudy (hba):

Lets see. find the Area A+B